OpenStudy (anonymous):

How many six-letter "words" can be formed from the letters A, B, C, D, E, F, if each letter can be used only once?

OpenStudy (anonymous):

6!

OpenStudy (anonymous):

these are the answers so i don't thinks its 6

OpenStudy (anonymous):

230,230 720 5040 890

OpenStudy (anonymous):

oh no i means \(6!=6\times 5\times 4\times 3\times 2\)

OpenStudy (anonymous):

Assuming you don't have to make real english words... You have six places to fill. So for the first space you have six choices. After you choose that one you have five for the next space. Then four for the next, etc. This is the same as 6! or 6x5x4x3x2x1

OpenStudy (anonymous):

(720)

OpenStudy (anonymous):

that comfused me dearly i was trying to make words and it lost me